Find all zeros of the polynomial equation
x3-2x2+10x+136=0

Answer by stanbon(57250)

Find all zeros of the polynomial equation
x^3-2x^2+10x+136=0
----------
I graphed it and found x=-4 is a zero.
Use synthetic division to find the other zeroes.
----------
-4)....1....-2....10....136
.........1....-6....34...|..0
Quotient: x^2 - 6x + 34
Use the quadratic formula to find remaining zeroes.
x = [6 +- sqrt(36 -4*1*34)]/2
x = [6 +- sqrt(-100)]/2
x = [6 +- 10i]/2
======================
Zeroes are x = -4, x = (3+5i), x = (3-5i)
=========================================
Cheers,
Stan H.